OpenStudy (anonymous):

The first four terms of a sequence are shown below. 1 4 9 16 The sequence continues. What is the tenth term of the sequence? 10 20 100 200

OpenStudy (anonymous):

it would be 100 right?

OpenStudy (texaschic101):

find the pattern... 1 + ? = 4 (3) 4 + ? = 9 (5) 9 + ? = 16 (7) it looks like the pattern is 3,5,7,9,11,13.... so the next number is 16 + 9 = 25 (5th term) next is 25 + 11 = 36 (6th term) 36 + 13 = 49 (7th term) 49 + 15 = 64 (8th term) 64 + 17 = 81 (9th term) 81 + 19 = 100 (10th term) yes...the 10th term is 100